OpenStudy (anonymous):

Kenny invested $5000 into a fund that is expected to grow by 3.5% per year. How long will it take the fund he invested in to be worth $10,000

OpenStudy (shiraz14):

This is a G.P. with a=5000, r=1.035 and Sn>=10,000. Find n using the formula of a G.P. as follows: Sn = [a(r^n -1)]/(r-1) Can you try it & post what your answer is in this thread?

OpenStudy (tkhunny):

No good, @shiraz14 This is only a single payment. \(5000(1.035)^{x} = 10000\) -- Solve for x. If you've heard of the "Rule of 72's", you can estimate rather well, 72/3.5 = 20.6 years

OpenStudy (shiraz14):

@tkhunny : Yes, you are correct - I must have been dreaming :) (I was thinking of compounded interest with the same fixed value invested into the account) ...

OpenStudy (anonymous):

10000 = (5000)(1.035)^x 2 =(1.035)^x ln 2 = x ln (1.035) x = 20.15 years, very close to the Rule of 72 value